OpenStudy (anonymous):

find both partial derivative z=sin(x+2y)

sam (.sam.):

wrt?

OpenStudy (anonymous):

\[z _{x}=\cos (x+2y)\] \[z _{y}=2\cos (x+2y)\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

wait. there's no symbol for the partial derivative here..

OpenStudy (anonymous):

its ok if there is no symbol of partial

OpenStudy (anonymous):

i already know how to solve this :)

OpenStudy (anonymous):

oh alright. had a hard time typing haha

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

i have a question :)

OpenStudy (anonymous):

Yeah?

OpenStudy (anonymous):

what is the derivative of cosh xy^2

OpenStudy (anonymous):

wrt to what?

OpenStudy (anonymous):

partial derivative

OpenStudy (anonymous):

respect to x .. y ..

OpenStudy (anonymous):

partial deriv wrt x is sinh(xy^2) partial wrt to y is (2y)sinh(xy^2)

OpenStudy (anonymous):

oops. partial wrt y = 2xy sinh(xy^2)

OpenStudy (anonymous):

not -sin?

OpenStudy (anonymous):

nope. there's a slight difference for hyperbolic functions :)

OpenStudy (anonymous):

a ok :)

OpenStudy (anonymous):

are you from the phils?

OpenStudy (anonymous):

yes why?

OpenStudy (anonymous):

napansin ko sa pag chat haha

OpenStudy (anonymous):

haha nose bleed ako filipino pala

OpenStudy (anonymous):

no capitalization, no period, "haha", and smileys

OpenStudy (anonymous):

ganun ba yun?

OpenStudy (anonymous):

oo. e ganun rin kase ako mag chat e

OpenStudy (anonymous):

a ok po san ka sa philippines? student?

OpenStudy (anonymous):

DLSU manila. u?

OpenStudy (anonymous):

wow!

OpenStudy (anonymous):

TIP manila

OpenStudy (anonymous):

ok malapit lang sa amin haha. hello i'm brett. kelangan ko na mag sign out :)

OpenStudy (anonymous):

thanks .. I'm airah :)